could you show me how to set this problem up. A collection of nickels, dimes and quarters totals $3.40. the total number of coins is 32, and there are twice as many dimes as quarters. how many of each are there.

Thank you!!

Answer by mananth(16949)

quarters------------dimes----------nickels
x------------------2x--------------32-3x
value = 340 cents total 32 coins ::: dimes & quarters = 3x
...
25x+10*2x+5(32-3x)=340
25x+20x+160-15x=340
30x=340-160
30x=180
/30
x=6 quarters
...
dimes = 2x= 12
nickels = 32-3x= 14
...
m.ananth@hotmail.ca
...